(-1, -2) (1, -1) What is the​ slope-intercept form in fraction form?

(-1, -2) (1, -1)

What is the​ slope-intercept form in fraction form?
Answer #1

Answer

We have (x1,y1) = (-1,-2) and (x2,y2) = (1,-1)

First, we need to find the slope of equation

slope = (y2-y1)/(x2-x1) = (-1-(-2))/(1-(-1)) = 1/2 ....................we used -(-a) =a identity

Now, using the equaton y =mx + b

setting (x,y) =(1,-1) and m = 1/2

we get

-1 = (1/2)*(1) + b

subtracting 1/2 on each side, we get

-1 - (1/2) = (1/2) - (1/2) + b

this gives us

b = - 1 -(1/2) = (-2-1)/2 = -3/2

So, we get b = -3/2 and m = 1/2

setting these values in general equation, we get

y = (1/2)x + (-3/2)